I have the answer but I would like to see how it is done.
An access ramp enters a building 1 m above ground and starts 3 m from the building. How long is the ramp.
1
Expert's answer
2012-05-24T08:00:59-0400
We'll need there Pythagorean theorem.
Denote lenght of the ramp by L. So we obtain
L^2=1^2+3^2=1+9=10
therefore L=sqrt(10)=approximately 3,16 m